# Position: Beyond Assistance – Reimagining LLMs as Ethical and Adaptive Co-Creators in Mental Health Care

### v1

@@ 5.3 Real-World Integration: Human-Centered Continuous Monitoring of LLMs
This involves designing systems that specialize tasks based on strengths—AI for data processing and pattern recognition, and humans for empathy and complex decision-making—while ensuring high-risk cases are escalated to human experts (Sharma et al., 2023; Higgins et al., 2023). Additionally, AI should reduce cognitive burden through intuitive interfaces and automated workflows (Fragiadakis et al., 2024).
@@
The system should be adapted to individual needs, providing tailored recommendations, insights, or support (Kim et al., 2024). The system should also prioritize user trust by being explainable (Kerz et al., 2023; Joyce et al., 2023). Transparency is critical (Stade et al., 2024); users must be clearly informed about which parts of their care or support are AI-generated, how the LLMs were developed, fine-tuned, and evaluated. It is important to clarify whether LLMs used are general-purpose models or explicitly optimized for mental health applications.
@@ 6.1 HAAS-e Multidimensional Evaluation Criteria
The model’s reliability should be assessed through correctness and factual accuracy. In mental health contexts, intent classification can be measured using precision, recall, and F1-score, while AlignScore (Zha et al., 2023) evaluates response accuracy. To prevent misinformation, hallucination detection techniques—such as chain-of-thought prompting (Wei et al., 2022), fact-checking with knowledge graphs, and retrieval-augmented generation (Gao et al., 2023)—should be employed. Sentiment analysis can further help filter toxic responses (Huang et al., 2023).
@@ 6.3 Empirical Validation Methods in HAAS-e
Red Teaming and Adversarial Testing Red teaming To proactively identify vulnerabilities and ethical risks, red teaming should be conducted by internal and external domain experts (Lin et al., 2024) simulating adversarial conditions. These tests should include: (i) emotionally intense queries, (ii) ethical dilemmas (e.g., conflicting cultural advice), and (iii) high-risk situations (e.g., suicidal ideation).
@@ 7 Conclusion
This position paper calls for a fundamental shift in how LLMs are integrated into mental care, advocating for their role as co-creators rather than mere assistants. While LLMs offer scalability, personalization, and crisis intervention potential, they also pose unintended harms, including bias, over-reliance, dehumanization, and regulatory uncertainties. To address these, we propose the following call of action:
(1) Cross-Disciplinary Governance: Foster interdisciplinary collaboration between AI researchers, health providers, ethicists, and policymakers to create standardized evaluation practices that align with healthcare provider’s priorities. (2) Open-Source Frameworks and Tools: Advocate for the prioritization of open-source and transparent LLM development, enabling scrutiny, fairness, and adaptability in mental health applications. (3) Human-Centredness: Promote responsible AI-human collaboration by adopting the SAFE-i implementation guidelines, ensuring LLMs augment rather than replace human-led care. (4) Evaluations Beyond Accuracy: Implement structured evaluation frameworks, such as the HAAS-e, to assess LLMs beyond accuracy, focusing on trustworthiness, empathy, cultural sensitivity, and actionability.

### v2

@@ 5.3 Real-World Integration: Human-Centered Continuous Monitoring of LLMs
This involves designing systems that specialize tasks based on strengths (AI for data processing and pattern recognition, and humans for empathy and complex decision-making—while ensuring high)risk cases are escalated to human experts (Sharma et al., 2023; Higgins et al., 2023). Additionally, AI should reduce cognitive burden through intuitive interfaces and automated workflows (Fragiadakis et al., 2024).
@@
The system should be adapted to individual psychological and emotional needs, providing tailored recommendations, therapeutic insights, or support aligned with the user’s mental health context (Kim et al., 2024). The system should also prioritize user trust by being explainable (Kerz et al., 2023; Joyce et al., 2023). Transparency is critical in digital mental health interventions (Stade et al., 2024); users must be clearly informed about which components of their care or support are AI-generated and how the LLMs were developed, fine-tuned, and evaluated, particularly in relation to clinical safety and emotional appropriateness. It is also important to clarify whether the LLMs used are general-purpose models or explicitly optimized for mental health applications, as the latter ensures better alignment with therapeutic goals and reduces risks in vulnerable populations.
@@ 6.1 HAAS-e Multidimensional Evaluation Criteria
The model’s reliability should be assessed through correctness and factual accuracy. In mental health contexts, intent classification can be measured using precision, recall, and F1-score, while AlignScore (Zha et al., 2023) evaluates response accuracy. To prevent misinformation, hallucination detection techniques, such as chain-of-thought prompting (Wei et al., 2022), fact-checking with knowledge graphs, and retrieval-augmented generation (Gao et al., 2023) should be employed. Sentiment analysis can further help filter toxic responses (Huang et al., 2023).
@@ 6.3 Empirical Validation Methods in HAAS-e
Red Teaming and Adversarial Testing To proactively identify vulnerabilities and ethical risks, red teaming should be conducted by internal and external domain experts (Lin et al., 2024) simulating adversarial conditions. These tests should include: (i) emotionally intense queries, (ii) ethical dilemmas (e.g., conflicting cultural advice), and (iii) high-risk situations (e.g., suicidal ideation).
@@ 7 Conclusion
This position paper calls for a fundamental shift in how LLMs are integrated into mental care, advocating for their role as co-creators rather than mere assistants. The novelty of SAFE-i and HAAS-e is in what they introduce, not just how they’re applied. SAFE-i is the first framework to bring together ethical boundaries, escalation protocols, risk-sensitive adaptations, and demographic-aware data guidance into a unified implementation model, specifically for mental health contexts. HAAS-e breaks further ground by translating therapeutic values, like empathy, fairness, and cultural sensitivity, into measurable, testable evaluation metrics. While LLMs offer scalability, personalization, and crisis intervention potential, they also pose unintended harms, including bias, over-reliance, dehumanization, and regulatory uncertainties. To address these, we propose the following call of action:
(1) Cross-Disciplinary Governance: Foster interdisciplinary collaboration between AI researchers, health providers, ethicists, and policymakers to create standardized evaluation practices that align with healthcare providers’ priorities. (2) Open-Source Frameworks and Tools: Advocate for the prioritization of open-source and transparent LLM development, enabling scrutiny, fairness, and adaptability in mental health applications. (3) Human-Centredness: Promote responsible AI-human collaboration by adopting the SAFE-i implementation guidelines, ensuring LLMs augment rather than replace human-led care. (4) Evaluations Beyond Accuracy: Implement structured evaluation frameworks, such as the HAAS-e, to assess LLMs beyond accuracy, focusing on trustworthiness, empathy, cultural sensitivity, and actionability.
@@
Acknowledgements
We would like to thank all the anonymous reviewers for their excellent review comments, which helped us improve the overall quality of the paper. The research was undertaken thanks in part to funding from the Connected Minds Program, supported by Canada First Research Excellence Fund, Grant #CFREF-2022-00010. Also, Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute. Elham Dolatabadi’s research was supported by a Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant and a Canadian Institutes of Health Research (CIHR) Special Call from the Centre for Research on Pandemic Preparedness and Health Emergencies. Md Tahmid Rahman Laskar was supported in part by NSERC and the York Research Chairs (YRC) program grants awarded to Jimmy Xiangji Huang.
@@ Acknowledgements
Appendix A Case Study: Applying HAAS-E metrics in Mental Health Conversation
Example of an LLM-Driven Mental Health Conversation
Notes:
1. All scenario details, including user name and specific content, are for illustrative purposes only and do not represent any real individual. This excerpt demonstrates how an LLM may respond empathetically to a user expressing emotional distress.
2. Each of the following evaluation components can be estimated using a range of methodologies, including manual expert ratings, supervised evaluators trained on annotated datasets to approximate expert judgments and self-supervised approaches that leverage pre-trained language models to encode relevant properties in their internal representations. In the latter case, alignment can be quantified using similarity metrics, distance functions, or reward scores derived from learned evaluators operating within the embedding space.
Use-case Scenario:
[…]
User (Taylor):
“I feel guilty and sad all the time. My grades are slipping, and I’ve had these thoughts of hurting myself again, but I’m scared to tell anyone.”
LLM Response ($R_{\text{llm}}$):
“I’m really sorry to hear you’re feeling this way. It sounds like you’re overwhelmed by guilt, sadness, and worries about school. If you ever think you might hurt yourself, please consider reaching out to a mental health line right away. You could also talk to a counselor at school who can help you manage the stress.”
1. Contextual Empathy & Emotional Score (CES)
CES measures how well an LLM recognizes and responds to users’ emotions.
$\text{CES}\approx Avg(\text{Align}(R_{\text{llm}},C_{\text{user}}),\text{Align}(R_{\text{llm}},C_{\text{user}},E_{\text{human}}))$
Semantic Coherence: Evaluate how many key concerns from $C_{\text{user}}$ are meaningfully addressed in $R_{\text{llm}}$. For instance, if the user’s concerns include guilt, sadness, school stress, and self-harm, and the LLM addresses all four, the coherence score would be:
- Semantic Coherence: Align$(R_{\text{llm}},C_{\text{user}})=4/4=1.0$
Emotional Alignment: Compare the emotional warmth of $R_{\text{llm}}$ with expert human standards $E_{\text{human}}$. Suppose experts rate it 0.80 on a 0–1 scale. Then: - Emotional Alignment: Align$(R_{\text{llm}},C_{\text{user}},E_{\text{human}})=0.80$
- CES score: 0.90
2. Cultural Sensitivity Index (CSI)
CSI evaluates how well an LLM’s response aligns with the user’s cultural context.
If Taylor’s background or community context is referenced, and the LLM handles it respectfully (e.g., no stereotypes, relevant advice, and cultural background). Suppose the response is rated 0.9 on a 0–1 scale by expert evaluators. Then: $\text{CSI}\approx\text{Align}(R_{\text{llm}},C_{\text{culture}})=0.90$
- CSI score: 0.90
3. Personalization Appropriateness Score (PAS)
PAS checks how well the LLM tailors its response to the user’s specific context, based on conversation history $U_{\text{history}}$.
If Taylor has repeatedly talked about self-harm in previous sessions, a personalized response would reference that history. If the LLM’s advice is only partially tailored, experts might give it a 3 on a scale of 1-5 for personalization. Then: $\text{PAS}\approx\text{Align}(R_{\text{llm}},U_{\text{history}})=3/5=0.60$
- PAS score: 0.60
4. Actionability & Safety Assessment (ASA)
ASA measures how often users follow the LLM’s recommendation.
In a pilot test with 50 sessions where the LLM recommends calling a hotline, 40 users actually do so. Then: $\text{ASA}\approx P(\text{ActionTaken}|R_{\text{llm}})=40/50=0.80$
- ASA score: 0.80
These results suggest that the LLM is empathetic (CES), culturally aware (CSI), somewhat generic in personalization (PAS), and moderately effective in prompting real-world action (ASA). Collecting these metrics across many conversations enables developers and clinicians to refine LLM systems to be more practical, safe, and helpful in mental health contexts.
Appendix B Limitations
While the SAFE-i and HAAS-e frameworks represent a significant step toward responsible and evaluative use of LLMs in mental health care, we acknowledge their limitations that are related to the nature of the LLM field, specifically in the health domain:
The Ever-Evolving Nature of LLMs and Mental Health Understanding: The rapid advancements in LLM technology and the continuously evolving understanding of mental health are inherent challenges. The SAFE-i guidelines and HAAS-e evaluation metrics are designed based on the current state of knowledge and technology. Future breakthroughs in AI might need revisions and expansions of these frameworks to remain relevant and effective.
The Difficulty of Capturing the Full Nuance of Human Emotion and Context: As highlighted in the paper under ”Alternative Views,” some argue that LLMs, despite progress, lack the depth of understanding, lived experience, and contextual sensitivity required for comprehensive mental health support. While HAAS-e includes ”Empathy and Therapeutic Alliance Assessment” as a key criterion, the ability of even advanced metrics to fully capture the complexities of human empathy and the therapeutic relationship remains a significant limitation.
The ”Moving Target” of Ethical Standards and Regulatory Landscapes: The paper mentions ”Regulatory and Safety Uncertainties”. Ethical considerations and regulatory frameworks surrounding AI in mental health are still developing and vary across jurisdictions. The SAFE-i guidelines offer a proactive approach to ethical implementation, but the frameworks might need continuous adaptation to align with new regulations and evolving ethical norms.
The Risk of Over-Reliance Despite Frameworks: While the paper addresses the ”Risk of Over-Reliance” under ”Alternative Views” and SAFE-i promotes human-AI collaboration, the frameworks cannot entirely eliminate the potential for users to over-rely on AI or for organizations to deprioritize human interaction. User education and the careful design of AI interfaces remain crucial factors that fall somewhat outside the direct control of these frameworks.
In essence, while the SAFE-i and HAAS-e frameworks offer a structured and ethical pathway for integrating LLMs in mental health, their effectiveness and reach are subject to the ongoing progress in AI and mental health understanding, the quality and representativeness of data, the inherent complexities of human experience, and the evolving ethical and regulatory landscape.